2 IV22 (3 points) R. In the circuit shown above, note the directions given for the conventional currents as well as for the polarity of the batteries. Use the following numerical values What is I1 What is I2? What is I? What is the power dissipated in R1? What is the power dissipated in R2? What is the power dissipated in R3? What is the power delivered by V? What is the power delivered by ½? A. W. W. W.
Answer #1

from krichoffs junction rule from loop rule 73-314,R-1.R = 0 42 = 74-34 (1) 31+71, +131, =0 31-71,-131 71-13(1,+I.) 31--201-